What is the length of a side of a square whose diagonal measures 4.24 ft?
Answer by nerdybill(7384)


Draw a diagram!
From it, you'll see that the diagonal and two sides form a right triangle.
Let x = length of a side of the square
then
x^2 + x^2 = 4.24^2
2x^2 = 4.24^2
2x^2 = 17.9776
x^2 = 8.9888
x = 3 feet